OpenStudy (anonymous):

y=6x^2-24x+28 put into y=a(x-h)^2+k form

OpenStudy (anonymous):

y=6(x^2-4x+4)+4 y=6(x^2-4x+2^2)+4 y=6(x-2)^2 + 4

OpenStudy (anonymous):

is there a maximum and minimum?